# Training-free LLM Verification via Recycling Few-shot Examples

### v1

@@ 5 Conclusion
We propose ReFeri, a training-free framework for selecting promising output from LLM that reuses few-shot examples not only for generation but also for validation. In our experiments, ReFeri performs consistently effective in various LLMs and tasks, demonstrating strong adaptability to RL-tuned LLMs and open-ended personalization settings, and highlighting its generality beyond reasoning tasks. Our results suggest that ReFeri is a practical way to find the reliable LLM output with minimal human involvement, opening future directions to reconsider the broader utility of demonstrations.
Limitation and future works. Since the selection by ReFeri is determined by likelihoods produced by an estimation model, it does not explain why a response is incorrect, unlike PRMs, which offer step-level feedback, or LLM-as-judge, which can easily generate explanations by prompting. However, we believe that ReFeri can potentially provide a certain level of interpretability; for example, we visualize the token-level uncertainty of candidate responses and observe that it reveals potentially untrustworthy tokens (see Appendix C). This kind of token-level consideration not only provides the interpretability but also can improve the effectiveness of ReFeri, suggesting a future direction. On the other hand, performing multiple inferences with the estimation model to obtain likelihoods for different few-shot examples (Eq. 8) can be costly, although it is still more efficient than generating responses with a state-of-the-art LLM like GPT-4o. However, as shown in Table 4, smaller estimation models can be highly effective depending on the task and domain; consequently, adaptively using models of varying scales of the estimation model could be effective in solving this issue.
@@ A.1 Datasets
  • MMLU-Pro. MMLU-Pro [40] is an extension of the original MMLU benchmark [41], which evaluates broad knowledge and reasoning over 57 subjects using 14k 4-way multiple-choice questions. MMLU-Pro introduces 12k curated 10-way multiple-choice questions across 14 professional domains, increasing task difficulty and emphasizing complex, multi-step reasoning. Instead of using the full test set, we subsample 300 questions per subject (totaling 4,200) using random seed 42 and we will share the used indices at the code. For few-shot examples, we follow the format used in [40]. This benchmark is used to assess domain-specific and robust reasoning performance.
@@
  • GPQA. GPQA [45] is a graduate-level QA benchmark consisting of 448 expert-authored multiple-choice questions in domains such as physics, chemistry, and biology. Designed to be "Google-proof," it focuses on evaluating complex scientific reasoning that cannot be answered through simple retrieval. We evaluate on GPQA-Diamond, a curated subset of 198 especially difficult questions selected by the authors. Few-shot examples are taken directly from the official release [45]. This task measures deep scientific understanding.
@@
  • DROP. The DROP benchmark [44] contains 96k question-answer pairs requiring discrete reasoning over Wikipedia passages (e.g., numerical operations, counting, or date comparison). Answers may include spans, numbers, or dates. We evaluate on a 500-sample subset randomly selected from the dev set, and we will share the selected indices at the code. We use 3-shot examples from [50] and report both EM and F1 metrics following the official implementation. This benchmark evaluates models’ symbolic reasoning grounded in natural language passages.
@@
  • HotpotQA. HotpotQA [42] consists of 113k multi-hop QA pairs requiring reasoning over multiple Wikipedia documents. The model must retrieve at least two relevant passages and combine facts to answer each question. We follow the [43], which uses 500 samples from the dev set. Few-shot examples are taken from [49]. This task tests compositional reasoning and the ability to aggregate distributed information across documents.
@@
  • MuSR. MuSR [46] is a benchmark for multi-step reasoning over long-form narratives (800–1000 words), constructed via neuro-symbolic generation to embed logical dependencies into natural language. It includes structured tasks such as TeamAllocation (constraint-based planning) and ObjectPlacement (spatial consistency reasoning). We evaluate on the 256 TeamAllocation and 250 ObjectPlacement examples from the official release [46], using 3-shot prompts tailored to each task [27]. MuSR requires understanding of narrative flow, contextual logic, and physical feasibility. As demonstrated in [27], ICL plays a critical role in model performance on MuSR, and demonstrates the effectiveness of ReFeri.
  • AIME24. AIME24 [57] consists of 30 official problems from the 2024 AIME I and II exams, widely known for their difficulty in symbolic math reasoning. These integer-answer questions cover topics such as algebra, combinatorics, and number theory, and are commonly used to evaluate mathematical reasoning capabilities of RLLMs. We use all 30 problems as-is and apply the same 5-shot prompting setup as in MATH500, following the implementation of [47].
  • AMC23. AMC23 [58] consists of 40 problems and includes selected high-difficulty problems from the 2023 AMC 12A and 12B exams. Evaluation is performed using the same 5-shot prompting setup as MATH500, following the implementation of [47]. This benchmark provides additional resolution for evaluating symbolic reasoning below the AIME level.
  • LaMP. The LaMP benchmark [15] evaluates personalized text generation across multiple tasks. We focus on LaMP-4 (personalized news headline generation) and LaMP-5 (personalized scholarly title generation). In LaMP-4, the model generates headlines conditioned on article content and author profile, while LaMP-5 requires generating research paper titles conditioned on researcher profiles. We follow the official test splits and evaluate with ROUGE-L and ROUGE-1. Few-shot RAG is an example of 5 shots retrieved from user history via BM25. These tasks test personalization and style-aware generation in open-ended outputs.
@@ A.2 Baselines
Here, we provide the template used for our baseline, using MATH500 as a representative task among multiple benchmarks. (see list DD).
@@ A.3 Implementation
Resource details. To avoid out-of-memory, we used two NVIDIA H100 GPUs for evaluation with the LLaMA-3.1-70B-Instruct model, and one H100 for DeepSeek-R1-Distill-Qwen-7B. All other experiments were performed on a single A6000 GPU.
Response generation. We use lm-eval-harness [^7] https://github.com/EleutherAI/lm-evaluation-harness to generate responses from LLaMA-based models, with temperature set to 1.0 and 5 responses sampled per input. The prompt was written in chat template format and used vllm. [^8] https://huggingface.co/datasets/meta-llama/Llama-3.1-8B-Instruct-evals For GPT-family models, we use the official OpenAI API to generate completions under the same sampling configuration. The remaining settings follow the GPT API default settings. During evaluation, we report the average score across the 5 generations. All evaluations are conducted using our custom evaluation scripts to ensure consistent scoring and formatting across models.
Algorithm of ReFeri. In algorithm 1, we present the formal algorithm for ReFeri. Except Table 4, we generate multiple candidate responses $\{r_{1},...,r_{K}\}$ for each test query using Few-shot CoT, as it exhibit the better quality on the average (see Table 1).
@@
Input: estimation model $P$, test-query $\widetilde{q}$, $K$ candidate responses $\{r_{1},\dots,r_{K}\}$, $N$ few-shot examples $\mathbf{X}=\{\mathbf{x}_{i}\}^{N}_{i=1}$, replaced prompt $\widetilde{\mathbf{X}}_{i}$, hyperparmeter $\beta$
@@
     $S_{\tt{Forw}}\leftarrow$ Compute forward score with $r_{k}$ as label, using $P$ and context $(\widetilde{q},\mathbf{X})$ (Eq. (6))
     $S_{\tt{Dire}}\leftarrow$ Compute direct score with $r_{k}$ as label, using $P$ and only $\widetilde{q}$
     Initialize backward score $S_{\tt{Back}}=0$
     for $i=1$ to $N$ do
         Construct $\widetilde{\mathbf{X}}_{i}\leftarrow$ using a leave-one-out strategy (Eq. (7))
         $S_{\tt{Back}}\leftarrow S_{\tt{Back}}+$ backward score with $a_{i}$ as label, using $P$ and $\widetilde{\mathbf{X}}_{i}$ (Eq. (8))
     end for
     $S_{\tt{Comb}}\leftarrow S_{\tt{Forw}}-S_{\tt{Back}}$ (Eq. (9))
     $S_{\tt{Fin}}\leftarrow(1-\beta)\cdot S_{\tt{Dire}}+\beta\cdot S_{\tt{Comb}}$ (Eq. (10))
@@
Generalization to reasoning LLM. We evaluate DeepSeek-R1-Distill-Qwen-7B [59] on two challenging math benchmarks: AMC23 and AIME24. For both datasets, we adopt the same 5-shot prompting used in MATH500, following the implementation of [47], as no task-specific few-shot examples are readily available and most prior works tend to prefer zero-shot formats due to frequent performance degradation under few-shot prompting [8]. We perform 5 generations per example using nucleus sampling ($p=0.95$), and temperature $1$, running vLLM with two H100 GPUs, with a maximum input and output length of 32,768 tokens. Our implementation is adapted from the FuseAI framework. [^9] https://github.com/fanqiwan/FuseAI/tree/main/FuseO1-Preview/math_evaluation We report the average of the five responses in the case of Zero-shot CoT and Few-shot CoT. To apply ReFeri, we follow the same procedure as in all other experiments. When a zero-shot setting is required, we simply substitute the model’s zero-shot response $r_{k}$ for each few-shot response in the same pipeline.
Application to LLM personalization. We experiment with two configurations: (1) vanilla generation, where only the user query is used as input without any personalization; and (2) RAG-based personalization, where the prompt is augmented with both a user profile and top-$5$ similar examples retrieved from the user’s history using BM25 [15, 60]. The responses are generated using the GPT-4o-mini model via the OpenAI API with temperature set to 1.0.
Building on the outputs generated through above pipeline, we apply our ReFeri method to select the most likely response among the five candidates for each input, with $\beta\in\{0.5,1\}$. For the evaluation, the ROUGE score between the response and the gold reference is used following prior works. Here, we observe that $\beta=1$ consistently yields better performance than $\beta=0.5$, which indicates the ineffectiveness of direct estimation in this task. This selection strategy shows improved alignment with the gold answers and performs favorably compared to vanilla generation and RAG-based personalization (see Table 5).
@@ Appendix B More Quantitative Results
B.1 Full results with different estimation models
Table 6: Full results with different estimation models on MATH500 and GPQA
(a) MATH500 (b) GPQA
Estimation GPT-4o-mini GPT-4o LLaMA3.1-8B Avg Estimation GPT-4o-mini GPT-4o LLaMA3.1-8B Avg
LLaMA-3.2-1B 78.0 77.8 52.0 69.3 LLaMA-3.2-1B 44.9 50.0 34.3 43.1
Qwen-2.5-7B 78.8 79.0 52.2 70.0 Qwen-2.5-7B 41.4 50.5 33.8 41.9
LLaMA-3.1-8B 78.2 78.4 51.0 69.2 LLaMA-3.1-8B 42.4 51.5 34.8 42.9
LLaMA-3.1-70B 77.8 78.0 53.4 69.7 LLaMA-3.1-70B 42.9 53.5 35.4 43.9
Table 6 provides full results for all estimation model combinations of MATH500 and GPQA. This complements the average performance across different generation LLMs (GPT-4o-mini, GPT-4o, and LLaMA3.1-8B) shown in Figure 4. On both tasks, ReFeri shows consistent performance gains regardless of the estimation model used, emphasizing its robustness. There are some model-specific trends; for example, smaller models (LLaMA-3.2-1B) perform competitively on (relatively) simple tasks like MATH500, as discussed in Section 4.2.
B.2 Additional ablation
Here, we conduct the additional experiments to provide comprehensive ablation study for ReFeri. First, we examine the effectiveness of the proposed prompt replacement (Eq. 7) for better estimation of backward score. To this end, we consider a simplified variant of our backward score, termed No replace, where each few-shot example $\mathbf{x}_{i}=(q_{i},a_{i})$ is evaluated in a one-shot manner using the test query $\widetilde{q}$ and the candidate response $r_{k}$ as additional context. Specifically, this variant modifies the backward score in Eq. (8) by replacing the leave-one-out prompt $\widetilde{\mathbf{X}}_{i}$ with a single pair $\mathbf{y}_{k}=(\widetilde{q},r_{k})$:
Next, we also examine the effectiveness of the direct likelihood term $\log P(\mathbf{y}_{k}):=\log P(r_{k}\mid\widetilde{q})$ when used alone (Direct). While this term is not required by the Bayes formulation, we initially hypothesized that it may serve as a useful prior when combined with other terms. However, as shown in Table 7, relying solely on the direct term yields lower accuracy (63.2%) than our scoring approach (63.6%). Remarkably, Direct slightly outperforms on MATH500 and GPQA, which are frequently encountered and handled domains (math and science), but it significantly underperforms on other domains which LLM should be adapted through in-context learning. This result supports our motivation of additional utilization of few-shot examples for the selection, to complement LLM’s limited intrinsic knowledge, especially for a new domain, through few-shot examples.
Table 7: Additional ablation study. Evaluation of scoring variants on responses by GPT-4o-mini on MATH500 and GPQA.
Methods MATH500 MMLU-pro GPQA DROP HotpotQA MuSR-op MuSR-ta Avg.
(Acc.) (Acc.) (Acc.) (EM / F1) (EM / F1) (Acc.) (Acc.)
No replace 78.4 65.0 42.4 78.4 / 84.1 36.2 / 48.0 60.2 82.4 63.3
Direct 78.4 64.9 42.9 79.2 / 84.9 35.2 / 47.1 59.8 82.0 63.2
ReFeri (Ours) 78.2 65.0 42.4 79.6 / 85.3 36.2 / 47.9 61.3 82.8 63.6
B.3 Additional comparison with few-shot prompting-based selection methods
@@
Models Methods MATH500 MMLU-pro GPQA DROP HotpotQA MuSR-op MuSR-ta Avg.
(Acc.) (Acc.) (Acc.) (EM / F1) (EM / F1) (Acc.) (Acc.)
GPT-4o-mini USC 78.6 62.5 42.4 78.8 / 85.8 36.6 / 48.2 59.8 76.4 62.2
USC-w/ Fewshot 79.4 62.9 38.9 77.8 / 85.5 36.2 / 48.3 59.0 77.2 61.6
LLM-as-Judge 79.0 64.7 39.9 78.6 / 85.6 36.2 / 48.4 58.6 77.2 62.0
ReFeri (Ours) 78.2 65.0 42.4 79.6 / 85.3 36.2 / 47.9 61.3 82.8 63.6
GPT-4o USC 79.8 72.1 50.5 82.0 / 90.2 45.8 / 60.4 73.4 87.6 70.2
USC-w/ Fewshot 80.2 72.2 47.5 82.6 / 90.7 46.2 / 60.7 71.1 89.6 69.9
LLM-as-Judge 80.8 76.6 51.0 82.4 / 90.2 46.6 / 61.3 72.3 89.2 71.3
ReFeri (Ours) 78.4 75.5 51.5 83.6 / 91.2 47.4 / 61.0 72.3 90.8 71.4
LLaMA-3.1-8B USC 49.6 35.6 28.8 69.6 / 75.8 24.4 / 32.5 52.3 67.2 46.8
USC-w/ Fewshot 47.8 36.5 28.3 69.0 / 75.3 25.2 / 32.3 53.9 70.0 47.2
LLM-as-Judge 46.0 44.1 21.2 67.6 / 74.0 23.4 / 31.2 55.1 66.4 46.2
ReFeri (Ours) 51.0 45.1 34.8 70.2 / 76.7 25.0 / 33.0 56.6 80.0 51.8
@@
MATH500 USC 87.8 9.4 1.0 0.4 1.4 0.0
USC-w/ Fewshot 89.2 8.6 0.8 0.2 1.2 0.0
LLM-as-Judge 91.0 4.6 1.0 2.4 1.0 0.0
MMLU-Pro USC 73.2 22.2 1.1 1.4 2.1 0.3
USC-w/ Fewshot 73.7 20.9 1.6 1.2 2.5 0.3
LLM-as-Judge 68.2 12.6 8.3 4.7 4.7 1.6
GPQA USC 66.2 29.3 0.5 0.0 4.0 0.0
USC-w/ Fewshot 65.7 25.3 3.0 1.0 5.1 0.0
LLM-as-Judge 37.4 26.3 16.7 10.6 5.1 4.0
DROP USC 81.8 15.4 1.2 0.6 1.0 0.0
USC-w/ Fewshot 82.8 13.8 1.2 0.8 1.4 0.0
LLM-as-Judge 85.0 9.2 2.2 2.8 0.8 0.0
HotpotQA USC 52.6 33.6 7.4 3.8 2.6 0.0
USC-w/ Fewshot 65.4 23.0 5.4 3.2 3.0 0.0
LLM-as-Judge 62.2 16.6 8.6 8.0 4.6 0.0
MuSR-op USC 91.4 7.42 0.4 0.4 0.4 0.0
USC-w/ Fewshot 92.6 7.42 0.0 0.0 0.0 0.0
LLM-as-Judge 85.2 7.81 3.9 1.2 2.0 0.0
MuSR-ta USC 55.6 31.2 2.4 1.2 0.0 9.6
USC-w/ Fewshot 60.8 33.2 6.0 0.0 0.0 0.0
LLM-as-Judge 34.4 36.4 272 0.0 2.0 0.0
Among the multiple answer selection methods, the simplest and most accessible approach (e.g., learning overhead, domain specificity, etc.) is arguably LLM-as-Judge [22, 23]. It uses the LLM itself to score and select answers via in-context learning without any additional training or external verifiers. In particular, the addition of few-shot examples to LLM-as-Judge might be most closely aligned with the core motivation of ReFeri, which is to use demonstrations not only for generation but also for validation. Therefore, in this section, we compare ReFeri and (1) the original USC [22], (2) USC with few-shot (our adaptation), and (3) LLM-as-Judge with few-shot created with our optimized prompt (see list D and D).

### v2

@@ 5 Conclusion
We propose ReFeri, a training-free framework to find promising LLM output by reusing few-shot data not only for generation but also for verification. In experiments, ReFeri performs consistently effective in various LLMs and tasks, demonstrating robustness across few-shot data and prompt variations. It suggests that ReFeri is a practical way to find the reliable LLM output with minimal human involvement, opening future directions to reconsider the broader utility of few-shot examples.
Limitation and future works. Since the selection by ReFeri is determined by likelihoods produced by an estimation model, it does not explain why a response is incorrect, unlike PRMs, which offer step-level feedback, or LLM-as-judge, which can easily generate explanations by prompting. However, we believe that ReFeri can potentially provide a certain level of interpretability; for example, we visualize the token-level uncertainty of candidate responses and observe that it reveals potentially untrustworthy tokens (see Appendix D). This kind of token-level consideration not only provides the interpretability but also can improve the effectiveness of ReFeri, suggesting a future direction.
Ethics Statement
ReFeri provides a training-free method for selecting promising outputs from LLMs. This makes it particularly valuable in scenarios where labeled data is scarce or where model fine-tuning is impractical such as limited access to data, or applications in emerging domains where predefined labels are unavailable. In addition, ReFeri reduces the barrier to deploying LLMs in real-world settings without additional supervision. This may contribute to broader and more efficient adoption of LLMs in resource-constrained environments. All datasets used are public and widely adopted.
Reproducibility Statement
For reproducibility, we provide detailed prompts, datasets, and experimental setups in Appendix A. In Section 3 and Appendix B, we report extensive experiments that demonstrate the robustness of our approach. In addition, we will release our code to ensure transparency and facilitate further research.
@@ A.1 Datasets
  • MMLU-Pro. MMLU-Pro (Wang et al., 2024b) is an extension of the original MMLU benchmark (Hendrycks et al., 2021a), which evaluates broad knowledge and reasoning over 57 subjects using 14k 4-way multiple-choice questions. MMLU-Pro introduces 12k curated 10-way multiple-choice questions across 14 professional domains, increasing task difficulty and emphasizing complex, multi-step reasoning. Instead of using the full test set, we subsample 300 questions per subject (totaling 4,200) using random seed 42 and we will share the used indices at the code. For few-shot examples, we follow the format used in (Wang et al., 2024b). This benchmark is used to assess domain-specific and robust reasoning performance.
@@
  • GPQA. GPQA (Rein et al., 2024) is a graduate-level QA benchmark consisting of 448 expert-authored multiple-choice questions in domains such as physics, chemistry, and biology. Designed to be "Google-proof," it focuses on evaluating complex scientific reasoning that cannot be answered through simple retrieval. We evaluate on GPQA-Diamond, a curated subset of 198 especially difficult questions selected by the authors. Few-shot examples are taken directly from the official release (Rein et al., 2024). This task measures deep scientific understanding.
@@
  • DROP. The DROP benchmark (Dua et al., 2019) contains 96k question-answer pairs requiring discrete reasoning over Wikipedia passages (e.g., numerical operations, counting, or date comparison). Answers may include spans, numbers, or dates. We evaluate on a 500-sample subset randomly selected from the dev set, and we will share the selected indices at the code. We use 3-shot examples from (Zhou et al., 2022) and report both EM and F1 metrics following the official implementation. This benchmark evaluates models’ symbolic reasoning grounded in natural language passages.
@@
  • HotpotQA. HotpotQA (Yang et al., 2018) consists of 113k multi-hop QA pairs requiring reasoning over multiple Wikipedia documents. The model must retrieve at least two relevant passages and combine facts to answer each question. We follow the (Kim et al., 2024), which uses 500 samples from the dev set. Few-shot examples are taken from (Yao et al., 2023). This task tests compositional reasoning and the ability to aggregate distributed information across documents.
@@
  • MuSR. MuSR (Sprague et al., 2024) is a benchmark for multi-step reasoning over long-form narratives (800–1000 words), constructed via neuro-symbolic generation to embed logical dependencies into natural language. It includes structured tasks such as TeamAllocation (constraint-based planning) and ObjectPlacement (spatial consistency reasoning). We evaluate on the 256 TeamAllocation and 250 ObjectPlacement examples from the official release (Sprague et al., 2024), using 3-shot prompts tailored to each task (Sprague et al., 2025). MuSR requires understanding of narrative flow, contextual logic, and physical feasibility. As demonstrated in (Sprague et al., 2025), ICL plays a critical role in model performance on MuSR, and demonstrates the effectiveness of ReFeri.
@@ A.2 Baselines
Here, we provide the template used for our baseline, using MATH500 as a representative task among multiple benchmarks. (see list 511).
@@ A.3 Implementation
Resource details. To avoid out-of-memory, we used two NVIDIA H100 GPUs for evaluation with the LLaMA-3.1-70B-Instruct model. All other experiments were performed on a single A6000 GPU.
Response generation. We use lm-eval-harness [^6] https://github.com/EleutherAI/lm-evaluation-harness to generate responses from LLaMA-based models, with temperature set to 1.0 and 5 responses per input. The prompt was written in chat template format using vllm. [^7] https://huggingface.co/datasets/meta-llama/Llama-3.1-8B-Instruct-evals For GPT-family models, we use the official OpenAI API to generate completions under the same sampling configuration. The remaining settings follow the GPT API default settings. During evaluation, we report the average score across the 5 generations. All evaluations are conducted using our custom evaluation scripts to ensure consistent scoring and formatting across models.
Algorithm of ReFeri. In algorithm 1, we present the formal algorithm for ReFeri. We generate multiple candidate responses $\{r_{1},...,r_{K}\}$ for each test query using Few-shot CoT, as it exhibit the better quality on the average (see Table 1).
@@
Input: estimation model $P$, embedding model $E$, test-query $\widetilde{q}$, $K$ candidate responses $\{r_{1},\dots,r_{K}\}$, $N$ few-shot examples $\mathbf{X}=\{\mathbf{x}_{i}\}^{N}_{i=1}$, replaced prompt $\widetilde{\mathbf{X}}_{i}$
$i^{\dagger}=\arg\max_{i=1,\dots,N}\cos\big(E(\widetilde{q}),E(q_{i})\big)$
@@
  $S_{\tt{Forw}}\leftarrow$ Compute forward score with $r_{k}$ as label, using $P$ and context $(\widetilde{q},\mathbf{X})$ (Eq. 6)
  Construct $\widetilde{\mathbf{X}}_{i^{\dagger}}\leftarrow$ using a leave-one-out strategy (Eq. 7)
  $S_{\tt{Back}}\leftarrow$ backward score with $a_{i}$ as label, using $P$ and $\widetilde{\mathbf{X}}_{i^{\dagger}}$ (Eq. 10)
  $S_{\tt{Fin}}\leftarrow S_{\tt{Forw}}-S_{\tt{Back}}$ (Eq. 11)
@@ Appendix B More Quantitative Results
B.1 Additional comparison with few-shot prompting-based selection methods
@@
Models Methods MuSR-ta MuSR-op GPQA MATH500 DROP HotpotQA MMLU-PRO Avg.
Llama-3.1-8B-It USC 67.2 52.3 28.8 49.6 69.6 / 75.8 24.4 / 32.5 45.6 48.2
USC-w/Fewshot 70.0 53.9 28.3 47.8 69.0 / 75.3 25.2 / 32.3 45.1 48.5
LLM-as-Judge 65.2 55.1 21.2 46.0 67.7 / 74.0 23.4 / 31.2 44.1 46.1
ReFeri 79.6 57.8 35.4 51.2 69.4 / 75.7 25.0 / 33.2 45.1 51.9
GPT-4o-mini USC 74.4 60.9 46.0 77.8 76.8 / 83.8 35.0 / 47.2 63.7 62.1
USC-w/Fewshot 76.4 63.3 39.9 78.2 77.2 / 84.0 34.8 / 46.6 63.2 61.9
LLM-as-Judge 75.6 60.6 34.3 77.0 77.4 / 84.4 35.0 / 46.7 63.3 60.5
ReFeri 82.8 61.3 41.9 77.8 79.2 / 84.9 36.2 / 48.0 64.9 63.4
GPT-4o USC 85.2 71.1 47.0 77.4 82.2 / 90.2 45.6 / 59.7 74.5 69.0
USC-w/Fewshot 88.8 69.1 46.0 77.4 82.0 / 89.9 45.4 / 60.1 74.1 69.0
LLM-as-Judge 86.0 68.0 46.5 77.8 82.8 / 91.0 45.6 / 59.8 73.3 68.6
ReFeri 90.4 71.9 51.5 77.8 83.6 / 91.1 47.0 / 60.7 75.4 71.1
Among the multiple answer selection methods, the simplest and most accessible approach (e.g., learning overhead, domain specificity, etc.) is arguably LLM-as-Judge Chen et al. (2023); Zheng et al. (2023). It uses the LLM itself to score and select answers via in-context learning without any additional training or external verifiers. In particular, the addition of few-shot examples to LLM-as-Judge might be most closely aligned with the core motivation of ReFeri, which is to use demonstrations not only for generation but also for validation. Therefore, in this section, we compare ReFeri and (1) the original USC (Chen et al., 2023), (2) USC with few-shot (our adaptation), and (3) LLM-as-Judge with few-shot created with our optimized prompt (see list 12 and 13).
@@
MATH500 USC 90.2 3.8 1.0 2.8 2.2 0.0
USC-w/ Fewshot 81.0 12.6 0.6 3.2 2.6 0.0
LLM-as-Judge 51.4 2.2 8.0 12.2 25.4 0.8
MMLU-Pro USC 34.2 19.1 7.2 24.0 15.5 0.0
USC-w/ Fewshot 18.8 31.0 7.8 22.7 19.7 0.0
LLM-as-Judge 22.1 8.0 7.8 13.5 48.3 1.2
GPQA USC 21.7 15.2 13.1 23.7 26.3 0.0
USC-w/ Fewshot 19.7 17.7 10.6 23.7 28.3 0.0
LLM-as-Judge 30.8 9.1 10.1 7.6 42.4 0.0
DROP USC 73.8 21.8 1.8 1.2 1.4 0.0
USC-w/ Fewshot 78.2 16.6 3.0 1.0 1.2 0.0
LLM-as-Judge 68.8 9.6 4.6 4.4 12.4 0.2
HotpotQA USC 77.0 13.4 2.8 5.4 1.4 0.0
USC-w/ Fewshot 68.6 20.6 2.8 6.0 2.0 0.0
LLM-as-Judge 65.0 15.0 4.8 6.0 9.2 0.0
MuSR-op USC 51.6 18.0 11.7 7.4 11.3 0.0
USC-w/ Fewshot 36.7 40.2 10.2 6.2 6.6 0.0
LLM-as-Judge 26.9 21.5 15.2 20.7 15.2 0.4
MuSR-ta USC 34.0 3.2 0.8 9.6 9.6 42.8
USC-w/ Fewshot 46.4 35.2 2.4 3.6 12.4 0.0
LLM-as-Judge 27.6 2.8 0.4 1.2 50.4 17.6
As shown in Table LABEL:tab:prompt_based_result, ReFeri consistently achieves the best or second-best accuracy across all LLMs and benchmarks. Interestingly, we observe that adding few-shot demonstrations to USC often degrades performance (e.g., on GPQA and DROP with GPT-4o-mini and LLaMA-3.1-8B), which is likely due to the sensitivity of LLMs to prompt format and positional bias of the responses.
Table 7: Evaluation of USC ordering with GPT-4o-mini. Two random permutations (perm-A, perm-B) of the candidate order versus the original order.
Methods MuSR-ta MuSR-op GPQA MATH500 DROP HotpotQA MMLU-PRO Avg.
USC (perm-A) 75.6 59.8 41.9 78.0 76.6 / 83.5 35.6 / 47.2 63.2 61.5
USC (perm-B) 77.2 56.6 45.0 77.2 76.6 / 83.4 35.2 / 46.9 63.6 61.6
USC (default) 74.4 60.9 46.0 77.8 76.8 / 83.8 35.0 / 47.2 63.7 62.1